OpenStudy (anonymous):

Choose the equation of the line passing through the point (-1, 7) and perpendicular to y = -one fourthx - 5. Choose one answer. a. y = -4x - 11 b. y = -4x - 3 c. y = 4x + 11 d. y = 4x + 3

OpenStudy (anonymous):

@AbdullahM

OpenStudy (anonymous):

for this im not sure what the slope would be but i know to just plug the rest in after

OpenStudy (abdullahm):

The slopes of 2 perpendicular lines when multiplied would be -1.

OpenStudy (anonymous):

so -1/4 * 4/1

OpenStudy (abdullahm):

Lets say the slope for the line they want us to find is 'y', then: \(\sf\Large \frac{-1}{4} \times y = -1\)

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

4/1

OpenStudy (abdullahm):

correct, so the slope if 4, because 4/1 = 4

OpenStudy (abdullahm):

What answer choice does that eliminate?

OpenStudy (anonymous):

a and b

OpenStudy (anonymous):

now i just plug them in again like the last one?

OpenStudy (abdullahm):

exactly :)

OpenStudy (anonymous):

c?

OpenStudy (abdullahm):

Correct \(\ddot\smile\)

OpenStudy (anonymous):

@AbdullahM do you know anything about correlations?

OpenStudy (abdullahm):

Somethings about it, post your question let me see :)